OpenStudy (anonymous):

Find the x and y intercepts and graph the equation plotting the intercepts. Please show all of your work. -5x + 8y = -40

OpenStudy (cwtan):

x-intercept y=0 -5x=-40 x=8 x-intercept=(8,0) y-intercept x=0 8y=40 y=5 y-intercept=(5,0)

OpenStudy (anonymous):

-5x + 8y = -40 -5x+8(0)=-40 x=8 (8,0) -5x + 8y = -40 -5(0)+8y=-40 y=5 (0,5)

OpenStudy (anonymous):

@cwtan y intercept should be 0,5

OpenStudy (anonymous):

thank you both tons!! I don't get this stuff very well. I appreciate it

OpenStudy (anonymous):

to get x int. change y=0, to get y int. change x=0

OpenStudy (anonymous):

ohh! ok! thank you!

OpenStudy (cwtan):

Oops~

OpenStudy (anonymous):



OpenStudy (anonymous):

it is a line

OpenStudy (anonymous):

ohh I see what you mean there. This stuff is just confusing, especially coming to finding slopes and things. i appreciate all the help!